OpenStudy (anonymous):

HELP - I do not understand word problems. **Use two variables to setup & solve** Kayaking: A kayaking group with a guide travels 16 miles downstream, stops for a meal, and then travels 16miles upstream. The speed of the current re mains constant throughout the trip. Find the speed of the kayak in still water. **************************************************** It then gives a table where it states: Leave @ 10 am Stop for meal @ 12pm Return @ 1pm Finish @ 5pm Let x=speed of kayak in still water Let y=speed of current

OpenStudy (callisto):

Basically, we apply the formula \[\large speed = \frac{distnace}{time}\] Note that the distance travelled is the same for travelling downstream and upstream Consider travelling downstream Speed = x+y Time = 2 hours Distance travelled = 16miles = 2(x+y) Consider travelling downstream Speed = x-y Time = 4 hours Distance travelled = 16miles = 4(x-y) So, 16 = 2(x+y) => 8 = x+y -(1) 16 = 4(x-y) => 4 = x-y -(2) (1) + (2) 2x = 12 x = 6 miles/hour 4 = x-y 4 = 6 -y y = ...? Not sure if it is correct...

OpenStudy (anonymous):

thank you - i checked the work and it's correct I get it now